# Asymptotic spreading of predator-prey populations in a shifting environment

### v1

@@ 4 Lower bound on the spreading speed
By the comparison principle, discussed in Section 4.3, the Hamilton-Jacobi equation (28) has a unique viscosity solution.
@@
For any given case $(i,j)\in\{(1,a),(1,b),(1,c),(2,a),(2,b)\}$ as stated in Section 3, let $R$ be given by $R=R_{ij}$. The Hamilton-Jacobi equation (28) has a unique viscosity solution, $\hat{\rho}$, satisfying
@@
$\hat{\rho}(s)\equiv 0\quad\text{ for }0\leq s\leq\sup\{s^{\prime}\geq 0:\leavevmode\nobreak\ \hat{\rho}(s)=0\}.$ (31)
@@
For any given case $(i,j)\in\{(1,a),(1,b),(1,c),(2,a),(2,b)\}$ as stated in Section 3, let $R$ be given by $R=R_{ij}$ and let $\hat{\rho}$ be the unique solution of (28) as specified in Lemma 7. Then
@@
For example, in Case 1(a) ($r_{1}<r_{2}$ and $c_{1}<2\sqrt{r_{2}}$), we will show that
@@ 4.1 Proof of Lemma 5
$o(1)=\partial_{t}\psi-\epsilon\partial_{xx}\psi+|\partial_{x}\psi|^{2}\leq\partial_{t}w^{\epsilon}-\epsilon\partial_{xx}w^{\epsilon}+|\partial_{x}w^{\epsilon}|^{2}=u^{\epsilon}-F^{\epsilon}\leq u^{\epsilon}-\delta_{0}\quad\text{ at }(t,x)=(t_{\epsilon},x_{\epsilon}),$ (35)
where $\delta_{0}=\beta\inf_{s\in\mathbb{R}}a(s)-1>0$.
@@ 4.2 Proof of Lemma 6
We recall the definition of viscosity solutions of (28).
@@
We recall the definition of viscosity solution of Hamilton-Jacobi equations with discountinuous Hamiltonian from [37]. In the following let $R^{*}$ and $R_{*}$ be the upper and lower envelope of $R$, which is given by
@@
We first show $w^{*}(t,x)\geq 0$ for $(t,x)\in[0,\infty)\times\mathbb{R}$. Indeed, since $u(t,x)\leq\max\{r_{2},r_{1}\}$ for all $(t,x)\in[0,\infty)\times\mathbb{R}$, by the definition of $w^{*}$ we have $w^{\epsilon}\geq-\epsilon\log(\max\{r_{2},r_{1}\})$ for each $\epsilon>0$ and $(t,x)\in[0,\infty)\times\mathbb{R}$, and we may compute
$w^{*}(t,x)=\limsup_{\begin{subarray}{c}\epsilon\to 0\\ (t^{\prime},x^{\prime})\to(t,x)\end{subarray}}w^{\epsilon}(t^{\prime},x^{\prime})\geq 0\quad\text{ for }(t,x)\in[0,\infty)\times\mathbb{R}.$ (39)
@@
$\inf_{|x|<\sqrt{r}t}u(t,x)\geq\delta_{1}\quad\text{ for }t\geq t_{1},$
@@
$\sup_{\begin{subarray}{c}|x|<\sqrt{r}t\\ t\geq\epsilon t_{1}\end{subarray}}w^{\epsilon}(t,x)\leq-\epsilon\log\delta_{1}.$ (43)
@@
$\rho^{*}(s_{0})-s_{0}\phi^{\prime}(s_{0})+|\phi^{\prime}(s_{0})|^{2}+R(s_{0})\leq 0.$
First, we note that $w^{*}(t,x)=\rho^{*}(x/t)$ and that $w^{*}(t,x)-t\phi(x/t)-(t-1)^{2}\leq 0$ for all $(t,x)$ near $(1,s_{0})$, with equality holding only at $(t,x)=(1,s_{0})$. Define, in terms of $\phi$, a two variable test function
@@
$\phi(s_{0})-s_{0}\phi^{\prime}(s_{0})+|\phi^{\prime}(s_{0})|^{2}+R(s_{0})\leq 0.$
Thus, $\rho^{*}$ is a viscosity sub-solution of (28).
@@ 4.3 Proof of Lemmas 7 and 8
5 Solving for the spreading speed via explicity formulas for $\hat{\rho}$
For each of the cases 1(a)-(c), 2(a), and 2(b), we will propose an explicit formula for $\hat{\rho}$ in Subsection 5.1. Thanks to the uniqueness result in Lemma 7, it is enough to verify (separately for each of the cases) that the given expression defines a viscosity solution of (28). This will be done in Subsection 5.2.
@@ 5.1 Explicit formulas for $\hat{\rho}$
    $\hat{\rho}(s)=\begin{cases}\frac{s^{2}}{4}-r_{2}\quad&\text{for }s>c_{1}\\ \lambda s-(\lambda^{2}+r_{1})\quad&\text{for }(\lambda+\frac{r_{1}}{\lambda})<s\leq c_{1}\\ 0\quad&\text{for }0\leq s\leq\lambda+\frac{r_{1}}{\lambda},\end{cases}$ (53)
    where $\lambda=\frac{c_{1}}{2}-\sqrt{r_{2}-r_{1}}$.
  • Case 1(c): $r_{1}<r_{2}$ and $c_{1}>2(\sqrt{r_{1}}+\sqrt{r_{2}-r_{1}})$
    $\hat{\rho}(s)=\begin{cases}\frac{s^{2}}{4}-r_{2}\quad&\text{for }s>c_{1}\\ \lambda s-(\lambda^{2}+r_{1})\quad&\text{for }2\lambda<s\leq c_{1}\\ \frac{s^{2}}{4}-r_{1}\quad&\text{for }2\sqrt{r_{1}}<s\leq 2\lambda\\ 0\quad&\text{for }0\leq s\leq 2\sqrt{r_{1}}.\end{cases}$ (54)
  • Case 2(a): $r_{1}>r_{2}$ and $c_{1}<2\sqrt{r_{2}}$.
@@
  • Case 2(b): $r_{1}>r_{2}$ and $c_{1}\geq 2\sqrt{r_{1}}$.
@@ 5.2 $\hat{\rho}$ solves the HJE (28) in the viscosity sense
We prove that the above formulas determine the the unique viscosity solution guaranteed by Lemma 7. To do that, it is enough to show that, in each case, $\hat{\rho}$ as given above, satisfies (i) the Hamilton-Jacobi equation (28) in the viscosity sense, as well as (ii) the boundary condition (30) in the classical sense. In view of the explicit formulas, (ii) is obvious. Thus, it remains to verify that $\hat{\rho}$ is a viscosity solution of (28) in each case. Since the proof only differs slightly in each case, we consider only two representative cases 1(b) and 2(b) here, and omit the verification of the rest.
@@
$R(s)=R_{1b}(s)=\begin{cases}r_{2}&\text{for }s\geq c_{1},\\ r_{1}&\text{for }\lambda+\frac{r_{1}}{\lambda}\leq s<c_{1}\\ \underline{r}_{1}&\text{for }s<\lambda+\frac{r_{1}}{\lambda}.\end{cases}$
@@
$\hat{\rho}(s)=\begin{cases}\frac{s^{2}}{4}-r_{2}\quad&\text{for }s>c_{1}\\ \lambda s-(\lambda^{2}+r_{1})\quad&\text{for }(\lambda+\frac{r_{1}}{\lambda})<s\leq c_{1}\\ 0\quad&\text{for }0\leq s\leq\lambda+\frac{r_{1}}{\lambda},\end{cases}$ (58)
where $\lambda=\frac{c_{1}}{2}-\sqrt{r_{2}-r_{1}}$.
First, observe that $\hat{\rho}$ is continuous, thanks to our choice of $\lambda$.
Next, we show that $\hat{\rho}$ is a viscosity subsolution of (28). To this end, observe that $\hat{\rho}$ satisfies the equation (28) in the classical sense almost everywhere in $[0,\infty)$. (In fact, it satisfies the equation classically for $s\in\mathbb{R}\setminus\{c_{1},\lambda+r_{1}/\lambda\}$.) By the convexity of the Hamiltonian, we can apply [4, Proposition 5.1] to conclude that it is in fact a viscosity sub-solution of (28).
Next, we show that $\hat{\rho}$ is a viscosity super-solution of (28). Suppose $\hat{\rho}-\phi$ obtains a strict local minimum at $s_{0}\in[0,\infty)$ for some $\phi\in C^{1}$. Now, $\hat{\rho}$ is a classical solution of (28) for all $s\not\in\{\lambda+\frac{r_{1}}{\lambda},c_{1}\}$, so it automatically satisfies (28) in the viscosity sense. We need only consider $s_{0}\in\{\lambda+\frac{r_{1}}{\lambda},c_{1}\}$. Suppose $s_{0}=\lambda+\frac{r_{1}}{\lambda}$. Then $0\leq\phi^{\prime}(s_{0})\leq\lambda$. Therefore, $R(s_{0})=r_{1}$ and at the point $s=s_{0}$, it holds that
$\rho-s_{0}\phi^{\prime}+|\phi^{\prime}|^{2}+R^{*}$ $=-\big{(}\lambda+\frac{r_{1}}{\lambda}\big{)}\phi^{\prime}+|\phi^{\prime}|^{2}+r_{1}=(\phi^{\prime}-\lambda)(\phi^{\prime}-\frac{r_{1}}{\lambda})\geq 0,$
where the last inequality is a consequence of $\phi^{\prime}(s_{0})\leq\lambda<\frac{r_{1}}{\lambda}$ (which in turn follows from the choice of $\lambda$ and the condition (57)).
@@
$\rho-s_{0}\phi^{\prime}+|\phi^{\prime}|^{2}+R^{*}$ $=\left(\tfrac{c_{1}^{2}}{4}-r_{2}\right)-c_{1}\phi^{\prime}+|\phi^{\prime}|^{2}+r_{2}$
$=(\phi^{\prime}-\frac{c_{1}}{2})^{2}\geq 0.$
@@
Let us proceed with Case 2(b), where we fix $r_{1},r_{2},c_{1}$ satisfying
$r_{1}>r_{2},\quad\text{ and }\quad c_{1}\geq 2\sqrt{r_{1}}.$ (59)
@@
$R(s)=R_{2b}(s)=\begin{cases}r_{2}&\text{for }s\geq c_{1},\\ r_{1}&\text{for }2\sqrt{r_{1}}\leq s<c_{1}\\ \underline{r}_{1}&\text{for }s<2\sqrt{r_{1}}\end{cases}$
@@
Again, we first observe that $\hat{\rho}$, as given in (60), is continuous, and satisfies the equation (28) in the classical sense for $s\in[0,\infty)\setminus\{c_{1},2\sqrt{r_{1}}\}$ (note that it is in fact continuously differentiable in a neighborhood of $s=2\tilde{\lambda}$). It then follows again from [4, Proposition 5.1] that $\hat{\rho}$ is a viscosity sub-solution of (28).
Next, we verify that it is also a viscosity super-solution. Suppose $\hat{\rho}-\phi$ obtains a strict local minimum at $s_{0}\in[0,\infty)$ for some $\phi\in C^{1}$. Since $\hat{\rho}$ is a classical solution of (28) for all $s\not\in\{c_{1},2\sqrt{r_{1}}\}$, we need only consider $s_{0}\in c_{1}$ or $s_{0}=2\sqrt{r_{1}}$. Suppose $s_{0}=c_{1}$, then
@@
This verifies that $\hat{\rho}$ is a viscosity super-solution of (28). This completes the proof that the unique viscosity solution $\hat{\rho}$ as guaranteed by Lemma 7 is given by the explicit formula (60) for the first representative Case 2(b).
@@
Suppose either that (1) $r_{1}<r_{2}$, or that (2) $r_{1}>r_{2}$ and $c_{1}\not\in(2\sqrt{r_{1}},2\sqrt{r_{2}})$. Then there exists $\delta_{0}>0$ such that for each $\eta>0$,
@@
where $\delta^{\prime}=\beta\inf a-1>0$, thanks to (H1). Since $u$ has compactly supported initial data, it follows from standard theory that the spreading speed of $u$ is bounded from below by $2\sqrt{\delta^{\prime}}$, i.e.
@@
For given $c_{1},r_{1},r_{2}>0$ satisfying any of the cases 1(a)-(c) and 2(a)-(b), Lemma 11 says that the unique solution $\hat{\rho}$ guaranteed in Lemma 7 is given as in Subsection 5.1. If we define

### v2

@@ 4 Lower bound on the spreading speed
By the comparison principle, discussed in the Appendix, the Hamilton-Jacobi equation (24) has a unique viscosity solution.
@@
For any given case $(i,j)\in\{(1,a),(1,b),(1,c),(2,a),(2,b),(2,c)\}$ as stated in Section 3, let $R$ be given by $R=R_{ij}$. The Hamilton-Jacobi equation (24) has a unique viscosity solution, $\hat{\rho}$, satisfying
@@
$\hat{\rho}(s)\equiv 0\quad\text{ for }0\leq s\leq\sup\{s^{\prime}\geq 0:\leavevmode\nobreak\ \hat{\rho}(s^{\prime})=0\}.$ (27)
@@
For any given case $(i,j)\in\{(1,a),(1,b),(1,c),(2,a),(2,b),(2,c)\}$ as stated in Section 3, let $R$ be given by $R=R_{ij}$ and let $\hat{\rho}$ be the unique solution of (24) as specified in Lemma 7. Then
@@
For example, in Case 1(a) ($r_{1}<r_{2}$ and $c_{1}\leq 2\sqrt{r_{2}}$), we find that
@@ 4.1 Proof of Lemma 5
$o(1)=\partial_{t}\psi-\epsilon\partial_{xx}\psi+|\partial_{x}\psi|^{2}\leq\partial_{t}w^{\epsilon}-\epsilon\partial_{xx}w^{\epsilon}+|\partial_{x}w^{\epsilon}|^{2}=u^{\epsilon}-F^{\epsilon}\leq u^{\epsilon}-\delta_{0}$ (31)
at $(t,x)=(t_{\epsilon},x_{\epsilon})$, where $\delta_{0}=\beta\inf_{s\in\mathbb{R}}a(s)-1>0$.
@@ 4.2 Proof of Lemma 6
We recall the definition of viscosity solutions of (24). We recall the definition of viscosity solution of Hamilton-Jacobi equations with discountinuous Hamiltonian, following [6] and originating from [38] (see also [25, 41] for the definition involving variational inequalities).
@@
In the following let $R^{*}$ and $R_{*}$ be the upper and lower envelope of $R$, which is given by
@@
We first show $w^{*}(t,x)\geq 0$ for $(t,x)\in(0,\infty)\times\mathbb{R}$. Indeed, since $u(t,x)\leq\max\{r_{2},r_{1}\}$ for all $(t,x)\in[0,\infty)\times\mathbb{R}$, by the definition of $w^{\epsilon}$ we have $w^{\epsilon}\geq-\epsilon\log(\max\{r_{2},r_{1}\})$ for each $\epsilon>0$ and $(t,x)\in(0,\infty)\times\mathbb{R}$, and we may compute
$w^{*}(t,x)=\limsup_{\begin{subarray}{c}\epsilon\to 0\\ (t^{\prime},x^{\prime})\to(t,x)\end{subarray}}w^{\epsilon}(t^{\prime},x^{\prime})\geq 0\quad\text{ for }(t,x)\in(0,\infty)\times\mathbb{R}.$ (33)
@@
$\inf_{|x|<\sqrt{\underline{r}}t}u(t,x)\geq\delta_{1}\quad\text{ for }t\geq t_{1},$
@@
$\sup_{\begin{subarray}{c}|x|<\sqrt{\underline{r}}t\\ t\geq\epsilon t_{1}\end{subarray}}w^{\epsilon}(t,x)\leq-\epsilon\log\delta_{1}.$ (35)
@@
$\rho^{*}(s_{0})-s_{0}\phi^{\prime}(s_{0})+|\phi^{\prime}(s_{0})|^{2}+R_{*}(s_{0})\leq 0.$
(Note that $R_{*}(s_{0})=R(s_{0})$ by our definition of $R$ in (19) and (20).) First, we note that $w^{*}(t,x)=\rho^{*}(x/t)$ and that $w^{*}(t,x)-t\phi(x/t)-(t-1)^{2}\leq 0$ for all $(t,x)$ near $(1,s_{0})$, with equality holding only at $(t,x)=(1,s_{0})$. Define, in terms of $\phi$, a two variable test function
@@
$\phi(s_{0})-s_{0}\phi^{\prime}(s_{0})+|\phi^{\prime}(s_{0})|^{2}+R_{*}(s_{0})\leq 0.$
Thus, $\rho^{*}$ is a viscosity sub-solution of (24). ∎
@@ 4.3 Proof of Lemmas 7 and 8
5 Solving for the spreading speed via explicit formulas for $\hat{\rho}$
For each of the cases 1(a)-(c), 2(a)-2(c), we will propose an explicit formula for $\hat{\rho}$ in Subsection 5.1. Thanks to the uniqueness result in Lemma 7, it is enough to verify (separately for each of the cases) that the given expression defines a viscosity solution of (24). This will be done in Subsection 5.2.
@@ 5.1 Explicit formulas for $\hat{\rho}$
    $\hat{\rho}(s)=\begin{cases}\frac{s^{2}}{4}-r_{2}\quad&\text{for }s>c_{1}\\ \lambda^{*}s-\big{(}(\lambda^{*})^{2}+r_{1}\big{)}\quad&\text{for }(\lambda^{*}+\frac{r_{1}}{\lambda^{*}})<s\leq c_{1}\\ 0\quad&\text{for }0\leq s\leq\lambda^{*}+\frac{r_{1}}{\lambda^{*}},\end{cases}$ (43)
    where $\lambda^{*}=\frac{c_{1}}{2}-\sqrt{r_{2}-r_{1}}$.
  • Case 1(c): $r_{1}<r_{2}$ and $c_{1}\geq 2(\sqrt{r_{1}}+\sqrt{r_{2}-r_{1}})$
    $\hat{\rho}(s)=\begin{cases}\frac{s^{2}}{4}-r_{2}\quad&\text{for }s>c_{1}\\ \lambda^{*}s-\big{(}(\lambda^{*})^{2}+r_{1}\big{)}\quad&\text{for }2\lambda^{*}<s\leq c_{1}\\ \frac{s^{2}}{4}-r_{1}\quad&\text{for }2\sqrt{r_{1}}<s\leq 2\lambda^{*}\\ 0\quad&\text{for }0\leq s\leq 2\sqrt{r_{1}}.\end{cases}$ (44)
  • Case 2(a): $r_{1}>r_{2}$ and $c_{1}<2\sqrt{r_{2}}$
@@
  • Case 2(b): $r_{1}>r_{2}$ and $c_{1}=2\sqrt{r_{2}}$
  • Case 2(c): $r_{1}>r_{2}$ and $c_{1}>2\sqrt{r_{1}}$.
@@ 5.2 $\hat{\rho}$ solves the HJE (24) in the viscosity sense
We prove that the above formulas determine the the unique viscosity solution guaranteed by Lemma 7. To do that, it is enough to show that, in each case, $\hat{\rho}$ as given above, satisfies (i) the Hamilton-Jacobi equation (24) in the viscosity sense, as well as (ii) the boundary condition (26) in the classical sense. In view of the explicit formulas, (ii) is obvious. Thus, it remains to verify that $\hat{\rho}$ is a viscosity solution of (24) in each case. Since the proof only differs slightly in each case, we consider only two representative cases 1(b) and 2(c) here, and omit the verification of the rest.
@@
$R(s)=R_{1b}(s)=\begin{cases}r_{2}&\text{for }s>c_{1},\\ r_{1}&\text{for }\lambda^{*}+\frac{r_{1}}{\lambda^{*}}<s\leq c_{1}\\ \underline{r}_{1}&\text{for }s\leq\lambda^{*}+\frac{r_{1}}{\lambda^{*}}.\end{cases}$
@@
$\hat{\rho}(s)=\begin{cases}\frac{s^{2}}{4}-r_{2}\quad&\text{for }s>c_{1}\\ \lambda^{*}s-\big{(}(\lambda^{*})^{2}+r_{1}\big{)}\quad&\text{for }(\lambda^{*}+\frac{r_{1}}{\lambda^{*}})<s\leq c_{1}\\ 0\quad&\text{for }0\leq s\leq\lambda^{*}+\frac{r_{1}}{\lambda^{*}},\end{cases}$ (49)
where $\lambda^{*}=\frac{c_{1}}{2}-\sqrt{r_{2}-r_{1}}$.
First, observe that $\hat{\rho}$ is continuous, thanks to our choice of $\lambda^{*}$.
Next, we show that $\hat{\rho}$ is a viscosity subsolution of (24). To this end, observe that $\hat{\rho}$ satisfies the equation (24) in the classical sense almost everywhere in $[0,\infty)$. (In fact, it satisfies the equation classically for $s\in\mathbb{R}\setminus\{c_{1},\lambda^{*}+r_{1}/\lambda^{*}\}$.) By the convexity of the Hamiltonian, we can apply [5, Proposition 5.1] to conclude that it is in fact a viscosity sub-solution of (24).
Next, we show that $\hat{\rho}$ is a viscosity super-solution of (24). Suppose $\hat{\rho}-\phi$ obtains a strict local minimum at $s_{0}\in[0,\infty)$ for some $\phi\in C^{1}$. Now, $\hat{\rho}$ is a classical solution of (24) for all $s\not\in\{\lambda^{*}+\frac{r_{1}}{\lambda^{*}},c_{1}\}$, so it automatically satisfies (24) in the viscosity sense. We need only consider $s_{0}\in\{\lambda^{*}+\frac{r_{1}}{\lambda^{*}},c_{1}\}$. Suppose $s_{0}=\lambda^{*}+\frac{r_{1}}{\lambda^{*}}$. Then $0\leq\phi^{\prime}(s_{0})\leq\lambda^{*}$, and $R^{*}(s_{0})=\max\{\underline{r}_{1},r_{1}\}=r_{1}$. Therefore, at the point $s=s_{0}$, it holds that
$\rho-s_{0}\phi^{\prime}+|\phi^{\prime}|^{2}+R^{*}$ $=-\big{(}\lambda^{*}+\frac{r_{1}}{\lambda^{*}}\big{)}\phi^{\prime}+|\phi^{\prime}|^{2}+r_{1}=(\phi^{\prime}-\lambda^{*})(\phi^{\prime}-\frac{r_{1}}{\lambda^{*}})\geq 0,$
where the last inequality is a consequence of $\phi^{\prime}(s_{0})\leq\lambda^{*}<\frac{r_{1}}{\lambda^{*}}$ (which in turn follows from the choice of $\lambda^{*}$ and the condition (48)).
@@
$\rho-s_{0}\phi^{\prime}+|\phi^{\prime}|^{2}+R^{*}$ $=\left(\tfrac{c_{1}^{2}}{4}-r_{2}\right)-c_{1}\phi^{\prime}+|\phi^{\prime}|^{2}+r_{2}=(\phi^{\prime}-\frac{c_{1}}{2})^{2}\geq 0\quad\text{ at }s=s_{0}.$
@@
Let us proceed with Case 2(c), where we fix $r_{1},r_{2},c_{1}$ satisfying
$r_{1}>r_{2},\quad\text{ and }\quad c_{1}>2\sqrt{r_{1}}.$ (50)
@@
$R(s)=R_{2c}(s)=\begin{cases}r_{2}&\text{for }s\geq c_{1},\\ r_{1}&\text{for }2\sqrt{r_{1}}<s<c_{1}\\ \underline{r}_{1}&\text{for }s\leq 2\sqrt{r_{1}}\end{cases}$
@@
$\hat{\rho}(s)=\begin{cases}\frac{s^{2}}{4}-r_{2}\quad&\text{for }s>2\tilde{\lambda}\\ \tilde{\lambda}s-(\tilde{\lambda}^{2}+r_{2})&\text{for }c_{1}<s\leq 2\tilde{\lambda}\\ \frac{s^{2}}{4}-r_{1}\quad&\text{for }2\sqrt{r_{1}}<s\leq c_{1}\\ 0\quad&\text{for }0\leq s\leq 2\sqrt{r_{1}},\end{cases}$ (51)
@@
Again, we first observe that $\hat{\rho}$, as given in (51), is continuous, and satisfies the equation (24) in the classical sense for $s\in[0,\infty)\setminus\{c_{1},2\sqrt{r_{1}}\}$ (note that it is in fact continuously differentiable in a neighborhood of $s=2\tilde{\lambda}$). It then follows again from [5, Proposition 5.1] that $\hat{\rho}$ is a viscosity sub-solution of (24).
Next, we verify that it is also a viscosity super-solution. Suppose $\hat{\rho}-\phi$ obtains a strict local minimum at $s_{0}\in[0,\infty)$ for some $\phi\in C^{1}$. Since $\hat{\rho}$ is a classical solution of (24) for all $s\not\in\{c_{1},2\sqrt{r_{1}}\}$, we need only consider $s_{0}=c_{1}$ or $s_{0}=2\sqrt{r_{1}}$. Suppose $s_{0}=c_{1}$, then
@@
This verifies that $\hat{\rho}$ is a viscosity super-solution of (24). This completes the proof that the unique viscosity solution $\hat{\rho}$ as guaranteed by Lemma 7 is given by the explicit formula (51) for the first representative Case 2(c).
@@
Suppose either that (1) $r_{1}<r_{2}$, or that (2) $r_{1}>r_{2}$ and $c_{1}\not\in(2\sqrt{r_{1}},2\sqrt{r_{2}}]$. Then there exists $\delta_{0}>0$ such that for each $\eta>0$,
@@
where $\delta^{\prime}=\beta\inf_{s\in\mathbb{R}}a(s)-1>0$, thanks to (H1). Since $u$ has compactly supported initial data, it follows from standard theory that the spreading speed of $u$ is bounded from below by $2\sqrt{\delta^{\prime}}$, i.e.
@@
For given $c_{1},r_{1},r_{2}>0$ satisfying any of the cases 1(a)-(c) and 2(a)-(c), Lemma 11 says that the unique solution $\hat{\rho}$ guaranteed in Lemma 7 is given as in Subsection 5.1. If we define